# Gravitational odd-parity perturbation of a Horndeski hairy black hole: quasinormal mode and parameter constraint

### v2

@@ V Conclusions and discussions
In this paper we have investigated how a specific Horndeski hair differ the QNMs of the black hole from GR in the context of gravitational odd perturbation, and intuitively analyzed the varying of QNMs with the hairy parameter by scrutinizing the evolution of the odd perturbation in time domain. Moreover, as a preliminary attempt, we used the yielding QNMs data to model ringdown waveforms of the hairy black hole and discussed the detectability of the hairy parameter, which was estimated from the Fisher information matrix for a single parameter. Our key findings are summarized as follows,
  • The radial propagation speed of gravitation wave emitted by the remnant black hole in Horndeski gravity at the current setup is found to be $c^{2}_{r}=(1+\frac{3Q}{2r})^{-1}$, which is different from the vacuum light speed in GR. To avoid Laplacian instability, we obtained a constraint on the black hole hair as $Q>-\frac{2}{3}r_{+}$.
@@
  • Comparing to the Shwarzschild black hole in GR, we found that the odd perturbation in the specific Horndeski gravity is significantly influenced by the Horndeski hair. We mainly analyzed the influence of the hairy parameter on the QNMs frequencies. We found that in both frequency domain and time domain, the hairy black hole is dynamically stable under the odd perturbation. In particular, our results shew that when $\tilde{Q}$ becomes negative enough, the leading mode becomes $\ell>2$ mode instead of $\ell=2$ mode, and also in this case the lifetime of the odd perturbation becomes the shortest one among all considered perturbations including massless scalar, Dirac, electromagnetic field perturbations. We argued that these phenomena may attribute to the superluminal propagation of the odd perturbation near the BH horizon, but the deeper physics deserves further study.
@@
  • In the context of parameter estimation by the ringdown waveform, we focused on the single parameter estimation of the Horndeski hair. Our results shew that the error bound of the mixed modes appeared in the average of the sum of corresponding single-mode errors. The ringdown observation would have a better constraint on the hairy parameter when the waveforms carry the larger $\ell$ QNMs with smaller node $n$. The Horndeski black hole with negative $\tilde{Q}$ was found to have a better constraint comparing to the cases with positive hairy charge. Furthermore, FIG.5-FIG.7 indicate that the rescaled measurement error of the hairy parameter is of order $\mathscr{O}(1-10)$. It means that the order of magnitude of the error on $\tilde{Q}$ for different GW detectors is around $\mathscr{O}(1-10)/\rho$ where $\rho$ is the Ringdown SNR for the corresponding detector. It is also worth noting that toward the end of this work, we became aware of a recent paper Sirera and Noller (2024) in which the authors forecast how GW experiments could constrain the scalar parameter of hairy black hole using only the fundamental $\ell=2$ ringdown mode extracted by the WKB method in scalar-tensor theory.
@@
Based on our present findings on the QNMs of gravitational odd perturbation and the error estimation, a series of further work can be proceeded to make the studies more physically realistic. Firstly, as an extension of the current work in the ringdown observation, a more comprehensive study on the parameter estimations, considering multi-parameters in the calculations of Fisher information matrix Berti et al. (2006a, b, 2009), will be a natural next step. We believe the covariance of general parameters (for instance, spins, masses, redshift, etc.) will produce more information about the detectability of a hairy black hole. Secondly, it is interesting to fit the ringdown wave form more realistically by imposing the non-linear modification Qiu et al. (2024) on the overtone model or considering the non-linear modes Khera et al. (2023) in constructing the wave forms. Thirdly, other polarizations of gravitational wave modes may carry more interesting property on the observation of black hole spectroscopy (see for example Barbeau et al. (2007); Rakhmanov (2005); Nishizawa et al. (2009); Gong and Hou (2018)), and the effect of Horndeski hair on the GWs in EMRI system could be complementary to our current study. Those are worth to be studied for future work. The last but not the least, the extension of these studies to the gravitational even perturbation in the Horndeski gravity would be another challenging direction. The studies on all these topics could be significant for our understanding on the no-hair theorem and its testing via GW experiments.

### v3

@@ V Conclusions and discussions
In this paper we have investigated how a specific Horndeski hair differ the QNMs of the black hole from GR in the context of gravitational odd perturbation, and intuitively analyzed the varying of QNMs with the hair parameter by scrutinizing the evolution of the odd perturbation in the time domain. Moreover, as a preliminary attempt, we used the yielding QNM data to model ringdown waveforms of the hairy black hole and discussed the detectability of the hair parameter, which was estimated from the Fisher information matrix for a single parameter. Our key findings are summarized as follows:
  • Proper choices of auxiliary field and tortoise coordinate can refine the modified Regge-Wheeler equation into a standard one, such that the superluminal propagation problem in the odd perturbation can be mathematically cured. Moreover, the refined equation has the same QNM spectrum as that for the original modified Regge-Wheeler equation.
@@
  • Comparing to the Shwarzschild black hole in GR, we found that the odd perturbation in the specific Horndeski gravity is significantly influenced by the Horndeski hair. We mainly analyzed the influence of the hair parameter on the QNMs frequencies. We found that in both frequency domain and time domain, the hairy black hole is dynamically stable under the odd perturbation. In particular, our results shew that when $\tilde{Q}$ becomes negative enough, the leading mode becomes $\ell>2$ mode instead of $\ell=2$ mode, and also in this case the lifetime of the odd perturbation becomes the shortest one among all considered perturbations including massless scalar, Dirac, electromagnetic field perturbations. We found that these abnormal effects can be traced to the mismatch between the null geodesic analysis and the gravitational odd perturbation in the eikonal limit. A deeper physical explanation deserves further study.
@@
  • In the context of parameter estimation by the ringdown waveform, we focused on the single parameter estimation of the Horndeski hair. Our results shew that the error bound of the mixed modes appeared in the average of the sum of corresponding single-mode errors. The ringdown observation would have a better constraint on the hair parameter when the waveforms carry the larger $\ell$ QNMs with smaller node $n$. The Horndeski black hole with negative $\tilde{Q}$ was found to have a better constraint comparing to the cases with positive hairy charge. Furthermore, FIG.6-FIG.8 indicate that the rescaled measurement error of the hair parameter is of order $\mathscr{O}(1-10)$. It means that the order of magnitude of the error on $\tilde{Q}$ for different GW detectors is around $\mathscr{O}(1-10)/\rho$ where $\rho$ is the Ringdown SNR for the corresponding detector. It is also worth noting that toward the end of this work, we became aware of a recent paper Sirera:2024ghv in which the authors forecast how GW experiments could constrain the scalar parameter of hairy black hole using only the fundamental $\ell=2$ ringdown mode extracted by the WKB method in scalar-tensor theory.
@@
    Based on our present findings on the QNMs of gravitational odd perturbation and the error estimation, a series of further work can be proceeded to make the studies more physically realistic. Firstly, as an extension of the current work in the ringdown observation, a more comprehensive study on the parameter estimations, considering multi-parameters in the calculations of Fisher information matrix Berti:2005ys ; Berti:2006hb ; Berti:2009kk , will be a natural next step. We believe the covariance of general parameters (for instance, spins, masses, redshift, etc.) will produce more information about the detectability of a hairy black hole. Secondly, it is interesting to fit the ringdown wave form more realistically by imposing the non-linear modification Qiu:2023lwo on the overtone model or considering the non-linear modes Khera:2023oyf in constructing the wave forms. Thirdly, other polarizations of gravitational waves modes may carry more interesting property on the observation of black hole spectroscopy (see for example Barbeau:2007qi ; Rakhmanov:2004eh ; Nishizawa:2009bf ; Gong:2017bru ), and the effect of Horndeski hair on the GWs in EMRI system could be complementary to our current study. Those are worth to be studied for future work. The last but not the least, the extension of these studies to the gravitational even perturbation in the Horndeski gravity would be another challenging direction, which is significantly required for a complete ringdown waveform construction. Though the even gravitational perturbation always couples with the scalar field in scalar tensor gravity, a method for the computation of QNMs frequencies in this coupled case was proposed in Langlois:2021aji and further applied in Roussille:2023sdr ; Roussille:2024xwr . The studies on all these topics could be significant for our understanding on the no-hair theorem and its testing via GW experiments.